OpenStudy (anonymous):

PLZ HELP!!!!! The formula r=2√5L can be used to approximate the speed r, in miles per hour, of a car that has left skid marks of length L, in feet. How far will a car skid at 20 mph? The car will skid ? feet.

OpenStudy (matt101):

\[2\sqrt{5L}=r\] \[\sqrt{5L}=r/2\] \[5L=r^2/4\] \[L=r^2/20\] Plugging in 20 mph for r, L=20 m.

OpenStudy (matt101):

sorry 20 feet

OpenStudy (anonymous):

OK? SO THE ANSWER IS...?

OpenStudy (matt101):

Just used to working in metric...

OpenStudy (anonymous):

?

OpenStudy (matt101):

The answer is 20 feet...

OpenStudy (anonymous):

THANK YOU